a. According to the Ministry of Health, the height of Guinean travellers who were quarantined in Tamale for the novel coronavirus were normally distributed about a mean of 160cm and a standard deviation of 8cm. Find the probability that a traveller selected at random has. Height between 148cm and 175cm, Height above 164cm, Height below 179cm. b.The average time taken for the nearest Covid-19 ambulance team to convey a patient showing symptoms to the nearest designated hospital follows the normal distribution about a mean of 60 seconds and a standard deviation of 8 seconds. If 200 patients were selected at random, how many of them could be taken to the hospital in. Less than 50 seconds, More than 64 seconds, The Shelf life of a particular dairy product is normally distributed with a mean of 12 days and a variance of 9 days. About what percentage of the products last between 13 and 16 days
1
Expert's answer
2020-06-08T19:23:00-0400

a)1) P(148<x<175)=P(x<175)-P(x<148)=

P(x<148)

z=(x-μ)/σ=(148-160)/8=-1.5

P(x<175)

z=(175-160)/8=1.875

P(148<x<175)=P(z<1.875)-P(z<-1.5)=0.9693−0.0668=0.2936

2)P(x>164)=1-P(x<164)

z=(164-160)/8=0.5

P(x>164)=1-P(z<0.5)=1-0.6915=0.3085

3)P(x<164)

z=(179-160)/8=2.37

P(x<164)=P(z<2.37)=0.9911

b) Mean =60 seconds

Sd=8 seconds

N=200 patients

i) the patients taken to hospital in less than 50 seconds

z=(50-60)/8=-1.25

P(z<-1.25) is 0.10565

The number of patients out of the 200 that will be taken to the hospital are

=0.10565*200≈21patients

ii) More than 64 seconds

z=(64-60)/8=0.5

P(z>0.5)=0.30854 from the standard normal tables

Out of the 200 randomly selected patients

=0.30854*200≈62patients will be taken to the hospital

с) mean=12

sd="\sqrt{ var}=3"

P(13<x<16)=P(x<16)-P(x<13)

P(x<13)

z=(x-μ)/σ=(13-12)/3=0.33

P(x<16)

z=(16-12)/3=1.33

P(13<x<16)=P(z<1.33)-P(z<0.33)=0.90824-0.62930=0.2789=27.89%